# TinySubNets: An efficient and low capacity continual learning strategy

### v1

@@ Abstract
In this subsection we describe more in detail an example of how TinySubNetworks manages memory capacity, as shown in Figure LABEL:fig:algorithm (with weight sharing, without replay memory, task 1, 2 and 3 share the weights, TSN-wr):
@@
A different view on capacity is shown in results in Table 5 (please note that results with replay memory include the memory occupation for the additional memory bank for each task - see Figure LABEL:fig:algorithm). Overall, we observe that post-pruning can reduce the memory footprint of the model. Its impact is particularly high with 5 Datasets (from $30.62\%$ to $24.68\%$). We can also observe that replay memory requires a consistent amount of capacity.
Results in Tables 6 and 7 allow us to highlight that the reduction in FLOPs provided by compression incurs a negligible accuracy loss. For instance, in the TinyImagenet scenario, 8 bit quantization allows to reduce from 4 775M FLOPs to 530.5M (or 450.9M with the pruning variant) at just 0.27 in accuracy drop. In Table 7 the accuracy after centroids quantization and activations quantization is described (in <4b, 4b>, <8b, 8b> and <16b, 16b> formats). It can be observed that in <8b, 8b> and <16b, 16b> configurations the accuracy is still at the same level. In case of <4b, 4b> there is a drop in accuracy. All these configurations can reduce number of FLOPs as it is shown in fig. 4. Note: Floating-point operations <flp, flp> have approximately 9× to 10× of the complexity of <8b, 8b> since there is 3× of the number of bits to represent the mantissas plus the remaining work required to handle exponents.
Forward and Backward Transfer are relevant metrics in continual learning. Results in Table 8 and Figures 2-5 show zero Backward Transfer for all scenarios, which is expected given the forget-free capabilities of the proposed method. As for Forward Transfer, values should be compared to the performance of a random classifier considering the number of classes in each dataset. Overall, we observe that Forward Transfer performance is close to or slightly above random. However, a proper evaluation of this metric is cumbersome in a task-incremental setting, given that our method requires knowledge of the task identifier to create masks and the specific sub-network for each task.
Heatmaps in Figures 2-5 provide additional disaggregated information, showing model performance on each task. It can be observed that the performance varies significantly depending on the task (e.g., on Imagenet-100, the performance varies from $68.9\%$ in task 5 to $83.8\%$ in task 7). The heatmaps further highlight the forget-free capabilities of the method, as shown by the performance on each task which is constant and preserved throughout the entire scenario.

### v2

@@ Acknowledgments
In this subsection we describe more in detail an example of how TinySubNetworks manages memory capacity, as shown in Figure 2 (with weight sharing, without replay memory, task 1, 2 and 3 share the weights, TSN-wr):
@@
A different view on capacity is shown in results in Table 8 (please note that results with replay memory include the memory occupation for the additional memory bank for each task - see Figure 2). Overall, we observe that post-pruning can reduce the memory footprint of the model. Its impact is particularly high with 5 Datasets (from $30.62\%$ to $24.68\%$). We can also observe that replay memory requires a consistent amount of capacity.
Results in Tables 9 and 10 allow us to highlight that the reduction in FLOPs provided by compression incurs a negligible accuracy loss. For instance, in the TinyImagenet scenario, 8 bit quantization allows to reduce from 4 775M FLOPs to 530.5M (or 450.9M with the pruning variant) at just 0.27 in accuracy drop. In Table 10 the accuracy after centroids quantization and activations quantization is described (in <4b, 4b>, <8b, 8b> and <16b, 16b> formats). It can be observed that in <8b, 8b> and <16b, 16b> configurations the accuracy is still at the same level. In case of <4b, 4b> there is a drop in accuracy. All these configurations can reduce number of FLOPs as it is shown in fig. 4. Note: Floating-point operations <flp, flp> have approximately 9× to 10× of the complexity of <8b, 8b> since there is 3× of the number of bits to represent the mantissas plus the remaining work required to handle exponents.
Forward and Backward Transfer are relevant metrics in continual learning. Results in Table 11 and Figures 6-9 show zero Backward Transfer for all scenarios, which is expected given the forget-free capabilities of the proposed method. As for Forward Transfer, values should be compared to the performance of a random classifier considering the number of classes in each dataset. Overall, we observe that Forward Transfer performance is close to or slightly above random. However, a proper evaluation of this metric is cumbersome in a task-incremental setting, given that our method requires knowledge of the task identifier to create masks and the specific sub-network for each task.
Heatmaps in Figures 6-9 provide additional disaggregated information, showing model performance on each task. It can be observed that the performance varies significantly depending on the task (e.g., on Imagenet-100, the performance varies from $68.9\%$ in task 5 to $83.8\%$ in task 7). The heatmaps further highlight the forget-free capabilities of the method, as shown by the performance on each task which is constant and preserved throughout the entire scenario.